OpenStudy (anonymous):

Complete the equation for the linear function whose graph contains the points (−1, −2) and (3, 10). y + ___ = ___ (x + 1)

OpenStudy (anonymous):

@Princer_Jones

OpenStudy (danjs):

y-(-2) = (10-(-2))/(3-(-1))* (x+1)

OpenStudy (anonymous):

im confused

OpenStudy (godlovesme):

Let's use the 1st ordered pair (-1,-2) Y-y1= m(x-x1) is called point-slope formula Do u know how to find slope? :)

OpenStudy (godlovesme):

@cjballer5323